Math 312.001: Introduction to Real Analysis

Stephen G. Simpson Pennsylvania State University 305 McAllister, 863-0775

                  simpson@math.psu.edu

January 21, 2009

This is an introductory course in real analysis. It is a transition course between lower-level mathematics courses (100 and 200 level) and higher-level mathematics courses (400 and 500 level).

The required textbook for this course is Elementary Analysis: The Theory of Calculus, by Kenneth A. Ross, Springer, X + 351 pages.

The web page for the course is

         http://www.math.psu.edu/simpson/courses/math312/.

We are meeting Mondays, Wednesdays, and Fridays, 2:30 PM to 3:20 PM, in 116 Osmond. Class attendance is mandatory. I am usually in my office Monday through Friday, 9:00 AM to 5:00 PM. Office hours are Monday and Wednesday 4:00–5:00 PM and by appointment, or you can drop in. Grades will be based on homework assignments (30 percent), two in-class midterm exams (40 percent), and a final exam (30 percent).

Academic Integrity Statement

Academic dishonesty is not limited to simply cheating on an exam or assignment. The following is quoted directly from the PSU Faculty Senate Policies for Students regarding academic integrity and academic dishonesty: “Academic integrity is the pursuit of scholarly activity free from fraud and deception and is an educational objective of this institution. Academic dishonesty includes, but is not limited to, cheating, plagiarizing, fabricating of information or citations, facilitating acts of academic dishonesty by others, having unauthorized possession of examinations, submitting work of another person or work previously used without informing the instructor, or tampering with the academic work of other students.” All University and Eberly College of Science policies regarding academic integrity/academic dishonesty apply to this course and the students enrolled in this course. Refer to the following URL for further details on the academic integrity policies of the Eberly College of Science:

         http://www.science.psu.edu/academic/Integrity/index.html

Each student in this course is expected to work entirely on her/his own while taking any exam, to complete assignments on her/his own effort without the assistance of others unless directed otherwise by the instructor, and to abide by University and Eberly College of Science policies about academic integrity and academic dishonesty. Academic dishonesty can result in assignment of “F” by the course instructors or “XF” by Judicial Affairs as the final grade for the student.